A random sample of n=36 observations is drawn from a population with a mean equal to25

and a standard deviation equal to 24

a.

Give the mean and standard deviation of the​ (repeated) sampling distribution x over bar x.

b.

Describe the shape of the sampling distribution of x over bar x. Does your answer depend on the sample​ size?

c.

Calculate the standard normal​ z-score corresponding to a value of x over bar x = 20.

d.

Calculate the standard normal​ z-score corresponding to a value of x over bar x=29.

a.

mu Subscript x over bar μ x =_____

sigma Subscript x over bar σ x =_______

b. Choose the correct answer below.

A. The shape is that of a normal distribution and depends on the sample size.

B. The shape is that of a uniform distribution and​ doesn't depend on the sample size.

C.The shape is that of a normal distribution and​ doesn't depend on the sample size.

D.The shape is that of a uniform distribution and depends on the sample size.

c. z=________

d. z=________
Answer #1

Please don't hesitate to give a "thumbs up" for the answer in case you're satisfied with the answer.

a.

Mean is equal to sample mean = 25
Stdev of this sample = stdev of population = 25/sqrt(n) = 25/sqrt(36) = 25/6

b.

It is approximately normal, yes it does depend on sample size

c.

for a x=20, z = (xbar-Mu)/(Stdev/sqrt(n)) = (20-25)/(24/sqrt(36)) = -1.25
d.

for a x=29, Z = (xbar-Mu)/(Stdev/sqrt(n)) = (29-25)/(24/sqrt(36)) = 1